LaTech Zill8e Ch4Sec4: Problem 3 Previous Problem Problem List Next Problem (1 point) Find the solution of y + 5y = 200 sinYou have to use Method of Undetermined Coefficients
Answer #1

Given, y + 5y = 200 sin (st) + 300 cos (5+) yo) = 8, yo = 4 with Il Auxillary equation is m?+ 5m = 0 - m (m + 5) = 0 L i mz0- 25 A sin (St) - 25 B Cos (5t) - + 25 A Cos ( 5t) - 25 B sin (5+) ___ = 200 sin (st) + 300 cos (5t) :: (-25A - 253) sin (5t)2. General solution - y = + Yp = 6 + (zēst + 2 sin (5+) -10 COS(SE) Now, we have y(0) = 8, y (o) = 4 .: 4 (0) = (y + lą ė + 2